# Spiking Transformers Need High Frequency Information

### v1

@@ 3.2.2 Patch Embedding
[…]
@@ 3.2.3 Token Mixing
In Transformers, lower layers typically require more high-frequency details, while higher layers benefit from more global information [28, 29]. Like biological vision, high-frequency details enable early stages to learn low-level features while progressively building local-to-global representations. Accordingly, we replace early-stage self-attention with DWC to preserve high frequencies essential for local feature learning. Given input embedding $\mathbf{Y}\in\mathbb{R}^{T\times C\times H\times W}$, the spiking DWC is defined as:
@@
where $\Omega(i)$ denotes the local neighborhood of position $i$, $w_{c,j}$ represents learnable convolution weights for channel $c$, and $X_{c},Z_{c}\in\mathbb{R}^{T\times H\times W}$ is the input and output slice for channel $c$. For the final stage, we implement token mixing via Spiking Self-Attention (SSA) [7]. The SSA computation follows:
@@ 3.3 Membrane Shortcut
Residual learning and shortcuts enable very deep networks to be trained effectively by providing identity shortcuts that preserve information flow and mitigate vanishing gradients [30]. For SNNs, a crucial consideration is maintaining compatibility with the spike-driven computing paradigm throughout all operations. As Fig. 5 shows, existing SNN shortcut implementations fall into three categories: (1) Vanilla Shortcut [31], (2) Pre-Spike Shortcut [32], and (3) Membrane Shortcut [33]. The Vanilla Shortcut scheme [31] directly connects spikes (binary) to membrane potentials (continuous), leading to a distribution mismatch that inherently violates the identity mapping principle. The Pre-Spike Shortcut [32] adds spike signals before neuron charging, resulting in summation values that range from 0 to 2. This disrupts the binary spike representation and spike-driven flow in SNNs.
In this work, we adopt the Membrane Shortcut [33] for its dual advantages: it preserves identity mapping by directly connecting membrane potentials while maintaining binary spike outputs that remain inherently compatible with spike-driven computation. Unlike Vanilla or Pre-Spike Shortcuts, this approach ensures both mathematical consistency with residual learning principles and seamless compatibility with spike-driven operations. We provide a detailed analysis of its impact on model performance and energy costs in Appendix A4.
@@ 4 Experiment
As shown in Figure 6, Max-Former restores high-frequency information by grafting the merits of frequency-enhancing operators: Max-Pooling and DWC. To empirically probe the importance of high frequency information in Spiking Transfomrers, we evaluate Max-Former through comprehensive experiments on static datasets (CIFAR-10 [34], CIFAR-100 [35] and ImageNet [36]) and neuromorphic datasets (CIFAR10-DVS [37], DVS128 Gesture [38]) , with architecture configurations detailed in Table 1. Experimental settings and energy analysis methods are detailed in Appendix A1 and A2.
Figure 6: Fourier spectrum of Spiking Neurons, Spiking Max-Pooling, Spiking Depth-Wise Convolution and Spiking Self-attention.
Table 1: Max-Former architecture configurations for different classification tasks. Notation DWC-$N$ denotes depth-wise convolution with kernel size $N\times N$. For block settings: CIFAR-10/100: 3 stages (1/1/2 blocks); ImageNet: 3 stages (1/3/7 blocks); Neuromorphic: 2 stages (1/1 blocks).
@@
Cifar10 [34]/100 [35] Embed-Orig Identity Embed-Max DWC-3 Embed-Max SSA
ImageNet [36] Embed-Orig DWD-7 Embed-Max DWC-5 Embed-Max SSA
Neuromorphic [37, 38] Embed-Max+ DWC-3 Embed-Max SSA
@@ 4.1 Results on CIFAR and Neuromorphic Datasets
Table 2: Performance Comparison on CIFAR10 [34], CIFAR100 [35], DVS128 [38] and CIFAR10-DVS [37]. Param.: Parameter (M); Acc.:Top-1 Accuracy (%); $T$: Simulation Timestep. *Models trained from scratch using identical configurations.
@@
ResNet-19 (ANN) [22] 12.63 1 94.97 12.63 1 75.35
Transformer (ANN) [22] 9.32 1 96.73 9.32 1 81.02
Spikformer [7] 9.32 4 95.51 9.32 4 78.21 2.57 16 98.3 2.57 16 80.9
S-Transformer [39] 10.28 4 95.60 10.28 4 78.40 2.57 16 99.3 2.57 16 80.0
SWformer [18] 7.51 4 96.10 7.51 4 79.30 - - - 2.05 16 83.9
QKFormer [22] 6.74 4 96.18 6.74 4 81.15 1.50 16 98.6 1.50 16 84.0
QKFormer* 6.74 4 96.84 6.74 4 81.57 1.50 16 98.6 1.50 16 82.3
@@
Max-Former and QKFormer share a similar hierarchical architecture, though QKFormer originally employs pre-spike shortcuts [22]. For a fair comparison, we additionally implemented QKFormer with the Membrane Shortcut (denoted as QKFormer* in the table) using identical training configurations. Max-Former still outperforms QKFormer* by 0.2% on CIFAR10 (97.04% vs. 96.84%) and by 1.08% on CIFAR100 (82.65% vs. 81.57%), while requiring slightly fewer parameters (6.57M vs. 6.74M). For neuromorphic datasets, Max-Former maintains this performance advantage. On DVS128, it achieves 98.6% accuracy, matching QKFormer* with membrane shortcut. On CIFAR10-DVS, Max-Former reaches 84.2% accuracy, exceeding QKFormer* (82.3%) by 1.9% and surpassing other spike-driven models like S-Transformer (80.0%) and SWformer (83.9%).
@@
ViT [40] ANN ViT-L/16 304.3 80.96 1 79.70
DeiT [41] ANN DeiT-B 86.6 80.50 1 81.80
PVT [42] ANN PVT-Large 61.4 45.08 1 81.70
MST [8] A2S Swin Transformer-T 28.50 512 78.51
Spikformer [7] SNN Spikformer-8-384 16.81 7.73 4 70.24
@@
QKFormer [22] SNN HST-10-384 16.47 15.13 4 78.80
@@ 4.2 Results on ImageNet Classification
Table 3 shows Max-Former’s performance on ImageNet classification, demonstrating its effectiveness for complex visual tasks. Max-Former-10-768 (T=4) achieves 82.39% accuracy (+7.58% over Spikformer) with 30% lower energy (14.87mJ vs 21.48mJ), despite using only lightweight for early-stage token mixing. It also outperforms the ANN-to-SNN MST model (78.51%) that requires 512 timesteps. Training/inference speed and memory usage are analyzed in Appendix A3.
Our analysis focuses on models using the Membrane Shortcut, which eliminates the energy-inefficient ternary spike transmissions ($\{0,1,2\}$ in Pre-Spike Shortcut) while maintaining full compatibility with standard neuromorphic hardware without additional hardware overhead (see Appendix A4). For fair comparison, we implemented a membrane potential variant of QKFormer (denoted QKFormer*), where QKFormer* shows 64% lower energy (5.52mJ vs 15.13mJ) in HST-10-384.
Max-Former-10-384 (16.23M, T=4) achieves 77.82% accuracy, outperforming QKFormer*-10-384 (16.47M, 76.48%), S-Transformer-8-768 (66.34M, 76.3%), and Meta-Spikformer (31.3M, 77.2%). For energy efficiency under identical settings, Max-Former-10-384 consumes 4.89mJ, significantly lower than QKFormer*-10-384(5.52mJ), S-Transformer(6.10mJ), and Meta-Spikformer (32.8mJ). Compared to conventional ANN models, Max-Former demonstrates concrete advantages in energy efficiency while maintaining competitive accuracy. Specifically, when compared to PVT-Large (a representative hierarchical ANN), Max-Former-10-768 (T=4) achieves comparable accuracy (82.34% vs. 81.70%) with 67% lower energy consumption (14.87mJ vs. 45.08mJ). These results confirm the importance of high-frequency information in Spiking Transformers: replacing energy-intensive self-attention with lightweight DWC in early stages actually produces better performance.
Table 4: Ablation of Patch-Embedding/ Token-Mixing Strategies on CIFAR100 and CIFAR10-DVS.
CIFAR100 CIFAR10-DVS
@@ 4.3 Ablation Study
Overall, Max-Former’s performance gains stem from effective high-frequency information propagation, independent of parameter count, which is evidenced by: (1) Max-pooling patch embedding (Embed-Max/Embed-Max+) consistently outperforms original versions despite similar parameter budgets. (2) Larger DWC kernels (DWC-5/DWC-7) increasing parameters but degrading accuracy (-0.63% on CIFAR100, -2.1% on CIFAR10-DVS vs Max-Former with DWC-3). See Appendix A4 for visualizations and A5 for limitations discussion.
@@ 5 Conclusion
This work challenges the prevailing assumption that binary activation constraints are the primary cause of Spiking Transformers’ performance gap. Through theoretical analysis and empirical validation, we demonstrate for the first time that spiking neurons inherently function as low-pass filters at the network level, resulting in the rapid attenuation of high-frequency components that critically degrade feature representation. To restore high-frequency signals, we design Max-Former with two key modifications: (1) additional Max-Pooling in patch embedding, and (2) lightweight Depth-Wise Convolution replacing early-stage self-attention for preserving local details. These frequency-enhancing designs enable Max-Former (63.99M parameters) to achieve 82.39% top-1 accuracy on ImageNet, representing a +7.58% improvement over Spikformer (74.81%, 66.34M parameters) while reducing energy consumption by 30% at a comparable model size. We believe this simple yet effective solution will motivate future research to explore the unique properties of spiking neural networks, beyond the established practices in standard deep learning paradigms.
Appendix A1 Analysis of High-Frequency Information in Spiking Transformers
To validate our approach, we conduct experiments across three static and two neuromorphic datasets. In this section, we first provide the detailed experimental setup used to obtain the results presented in our paper. We then perform additional analysis on the importance of high-frequency information in Spiking Transformers. The complete codebase and more comprehensive parameter settings will be made available upon acceptance.
@@
Resolution $224\times 224$ $32\times 32$ $32\times 32$ $128\times 128$
@@
A1.1 Experimental Details
Datasets: We evaluate Max-Former through comprehensive experiments on static datasets (CIFAR-10 [34], CIFAR-100 [35] and ImageNet [36]) and neuromorphic datasets (CIFAR10-DVS [37], DVS128 Gesture [38]). The training and inference pipeline are implemented in SpikingJelly [43].
Static Datasets: For static image classification, we evaluate on three standard benchmarks. ImageNet-1k [36] is one of the most widely used datasets in computer vision. It contains 1.28 million training, 50,000 validation, and 100,000 test images covering the common 1K classes. Both CIFAR-10 [34] and CIFAR-100 [35] include 50,000 training images and 10,000 testing images with 32$\times$32 resolution. The main difference between them is that CIFAR-10 has 10 categories for classification, while CIFAR-100 has 100 categories.
Neuromorphic Datasets: For event-based vision tasks, we evaluate on two standard benchmarks. CIFAR10-DVS [37] is an event-based version of the CIFAR-10 dataset, created by capturing moving image samples using the Dynamic Vision Sensor (DVS). It includes 10,000 event-based images (128$\times$128 pixels) spread across 10 classes, with 9,000 samples for training and 1,000 for testing. The DVS128 Gesture dataset [38] contains 1,342 event-based recordings of 11 different hand gesture types performed by 29 people under 3 different lighting conditions. Each gesture recording lasts about 6 seconds on average.
Hyper Parameters: Our training scheme mainly follows [22] and  [39]. Specifically, MixUp [44], CutMix [45] and RandAugment [46] are used for data augmentation. The models are trained using AdamW optimizer [47] with the weight decay of 0.05 for ImageNet-1K classification tasks and the weight decay of 0.06 for all other datasets. Label Smoothing [48] is set as 0.1. Detailed training hyperparameters are shown in Table 5. For our ImageNet experiments, we used 8 NVIDIA A30 GPUs to train most models. However, for the MaxFormer-10-512 (T=4) and MaxFormer-10-768 (T=4) models, we used 8 NVIDIA H20 GPUs instead. For the smaller datasets (CIFAR10, CIFAR100, DVS128 Gesture, and CIFAR10-DVS), we used a single A30 GPU for training.
A1.2 Impact of High-Frequency Information on Model Performance
Table 6: Patch embedding and token mixing schemes on CIFAR-100. Acc.:Top-1 Accuracy (%). DWC-$N$: spiking depth-wise convolution with kernel size $N\times N$. SSA: Spiking Self-Attention.

### v6

@@ 3.2.2 Patch Embedding
$\text{Embed-Orig~~~:}(\mathcal{G}_{1},\mathcal{G}_{2})=\text{(Embed, Embed)}$ (16)
$\text{Embed-Max~~~:}(\mathcal{G}_{1},\mathcal{G}_{2})=\text{(Max-Embed, Embed)}$ (17)
$\text{Embed-Max+ :}(\mathcal{G}_{1},\mathcal{G}_{2})=\text{(Max-Embed, Max-Embed)}$ (18)
where: $\text{Embed}\equiv\{\text{LIF - CONV - BN}\}$, and $\text{Max-Embed}\equiv\{\text{LIF - CONV - BN - MaxPool}\}$.
@@ 3.2.3 Token Mixing
In Transformers, lower layers typically require more high-frequency details, while higher layers benefit from more global information [raghu2021vision, si2022inception]. Like biological vision, high-frequency details enable early stages to learn low-level features while progressively building local-to-global representations. Accordingly, we replace early-stage self-attention with DWC to preserve high frequencies essential for local feature learning. Given input embedding $\mathbf{Y}\in\mathbb{R}^{T\times C\times H\times W}$, the spiking DWC is defined as:
@@
where $\Omega(i)$ denotes the local neighborhood of position $i$, $w_{c,j}$ represents learnable convolution weights for channel $c$, and $X_{c},Z_{c}\in\mathbb{R}^{T\times H\times W}$ is the input and output slice for channel $c$. For the final stage, we implement token mixing via Spiking Self-Attention (SSA) [zhou2022spikformer]. The SSA computation follows:
@@ 3.3 Membrane Shortcut
Residual learning and shortcuts enable very deep networks to be trained effectively by providing identity shortcuts that preserve information flow and mitigate vanishing gradients [he2016identity]. For SNNs, a crucial consideration is maintaining compatibility with the spike-driven computing paradigm throughout all operations. As Fig. 5 shows, existing SNN shortcut implementations fall into three categories: (1) Vanilla Shortcut [zheng2021going], (2) Pre-Spike Shortcut [fangDeepResidualLearning2021], and (3) Membrane Shortcut [hu2024advancing].
The Vanilla Shortcut scheme [zheng2021going] directly connects spikes (binary) to membrane potentials (continuous), leading to a distribution mismatch that inherently violates the identity mapping principle. The Pre-Spike Shortcut [fangDeepResidualLearning2021] adds spike signals before neuron charging, resulting in summation values that range from 0 to 2. This disrupts the binary spike representation and spike-driven flow in SNNs.
In this work, we adopt the Membrane Shortcut [hu2024advancing] for its dual advantages: it preserves identity mapping by directly connecting membrane potentials while maintaining binary spike outputs that remain inherently compatible with spike-driven computation. Unlike Vanilla or Pre-Spike Shortcuts, this approach ensures both mathematical consistency with residual learning principles and seamless compatibility with spike-driven operations. We provide a detailed analysis of its impact on model performance and energy costs in Appendix D.
@@ 4 Experiment
As shown in Figure 6, Max-Former restores high-frequency information by grafting the merits of frequency-enhancing operators: Max-Pool and DWC. To empirically probe the importance of high-frequency information in Spiking Transformers, we evaluate Max-Former through comprehensive experiments on static datasets (CIFAR-10 [lecunGradientbasedLearningApplied1998a], CIFAR-100 [krizhevskyLearningMultipleLayers2009a] and ImageNet [dengImagenetLargescaleHierarchical2009]) and neuromorphic datasets (CIFAR10-DVS [liCIFAR10DVSEventStreamDataset2017], DVS128 Gesture [amir2017low]), with architecture configurations detailed in Table 1. In addition, we design Max‑ResNet to further investigate the effect of high‑frequency restoration in convolutional architectures (model implementation is detailed in Appendix A.3). Experimental settings and energy analysis methods are detailed in Appendix A and B.
Figure 6: Fourier spectrum of Spiking Neurons, Spiking Max-Pool, Spiking Depth-Wise Convolution and Spiking Self-attention.
Table 1: Max-Former architecture configurations for different classification tasks. Notation DWC-NN denotes depth-wise convolution with kernel size $N\times N$. For block settings: CIFAR-10/100: 3 stages (1/1/2 blocks); ImageNet: 3 stages (1/3/7 blocks); Neuromorphic: 2 stages (1/1 blocks).
@@
Cifar10 [lecunGradientbasedLearningApplied1998a]/100 [krizhevskyLearningMultipleLayers2009a] Embed-Orig Identity Embed-Max DWC-3 Embed-Max SSA
ImageNet [dengImagenetLargescaleHierarchical2009] Embed-Orig DWD-7 Embed-Max DWC-5 Embed-Max SSA
Neuromorphic [liCIFAR10DVSEventStreamDataset2017, amir2017low] Embed-Max+ DWC-3 Embed-Max SSA
@@ 4.1 Results on CIFAR and Neuromorphic Datasets
Table 2: Performance Comparison on CIFAR10 [lecunGradientbasedLearningApplied1998a], CIFAR100 [krizhevskyLearningMultipleLayers2009a], DVS128 [amir2017low] and CIFAR10-DVS [liCIFAR10DVSEventStreamDataset2017]. Param.: Parameter (M); Acc.:Top-1 Accuracy (%); $T$: Simulation Timestep. *Models trained from scratch using identical configurations.
@@
ResNet-19 (ANN) [zhou2024qkformer] 12.63 1 94.97 12.63 1 75.35
Max-Former (ANN) 6.57 1 96.82 6.60 1 82.41
Spikformer [zhou2022spikformer] 9.32 4 95.51 9.32 4 78.21 2.57 16 98.3 2.57 16 80.9
S-Transformer [yao2023spike] 10.28 4 95.60 10.28 4 78.40 2.57 16 99.3 2.57 16 80.0
SWformer [fang2024spiking] 7.51 4 96.10 7.51 4 79.30 - - - 2.05 16 83.9
QKFormer [zhou2024qkformer] 6.74 4 96.18 6.74 4 81.15 1.50 16 98.6 1.50 16 84.0
MS-QKFormer* 6.74 4 96.84 6.74 4 81.57 1.50 16 98.6 1.50 16 82.3
@@
ViT [vaswani2017attention] ANN ViT-L/16 304.3 80.96 1 79.70
DeiT [touvron2021training] ANN DeiT-B 86.6 80.50 1 81.80
PVT [wang2021pyramid] ANN PVT-Large 61.4 45.08 1 81.70
MST [wang2023masked] A2S Swin Transformer-T 28.50 512 78.51
Spikformer [zhou2022spikformer] SNN Spikformer-8-384 16.81 7.73 4 70.24
@@
SWformer [fang2024spiking] SNN SWformer-8-512 27.6 5.08 4 75.43
QKFormer [zhou2024qkformer] SNN HST-10-384 16.47 15.13 4 78.80
@@
Max-Former and QKFormer share a similar hierarchical architecture, though QKFormer originally employs pre-spike shortcuts [zhou2024qkformer]. For a fair comparison, we additionally implemented QKFormer with the Membrane Shortcut (denoted as MS-QKFormer in the table) using identical training configurations. Max-Former still outperforms MS-QKFormer by 0.2% on CIFAR10 (97.04% vs. 96.84%) and by 1.08% on CIFAR100 (82.65% vs. 81.57%), while requiring slightly fewer parameters (6.57M vs. 6.74M). For neuromorphic datasets, Max-Former maintains this performance advantage. On DVS128, it achieves 98.6% accuracy, matching MS-QKFormer with the membrane shortcut. On CIFAR10-DVS, Max-Former reaches 84.2% accuracy, exceeding MS-QKFormer (82.3%) by 1.9% and surpassing other spike-driven models like S-Transformer (80.0%) and SWformer (83.9%).
@@ 4.2 Results on ImageNet Classification
Table 3 shows Max-Former’s performance on ImageNet classification, demonstrating its effectiveness for complex visual tasks. Max-Former-10-768 (T=4) achieves 82.39% accuracy (+7.58% over Spikformer) with 30% lower energy (14.87mJ vs 21.48mJ), despite using only lightweight for early-stage token mixing. It also outperforms the ANN-to-SNN MST model (78.51%) that requires 512 timesteps. Training/inference speed and memory usage are analyzed in Appendix C.
Our analysis focuses on models using the Membrane Shortcut, which eliminates the energy-inefficient ternary spike transmissions ($\{0,1,2\}$ in Pre-Spike Shortcut) while maintaining full compatibility with standard neuromorphic hardware without additional hardware overhead (see Appendix D). For fair comparison, we implemented a membrane potential variant of QKFormer (denoted MS-QKFormer), where MS-QKFormer shows 64% lower energy (5.52mJ vs 15.13mJ) in HST-10-384.
Max-Former-10-384 (16.23M, T=4) achieves 77.82% accuracy, outperforming MS-QKFormer-10-384 (16.47M, 76.48%), S-Transformer-8-768 (66.34M, 76.3%), and Meta-Spikformer (31.3M, 77.2%). For energy efficiency under identical settings, Max-Former-10-384 consumes 4.89mJ, significantly lower than MS-QKFormer-10-384(5.52mJ), S-Transformer(6.10mJ), and Meta-Spikformer (32.8mJ). Compared to conventional ANN models, Max-Former demonstrates concrete advantages in energy efficiency while maintaining competitive accuracy. Specifically, when compared to PVT-Large (a representative hierarchical ANN), Max-Former-10-768 (T=4) achieves comparable accuracy (82.34% vs. 81.70%) with 67% lower energy consumption (14.87mJ vs. 45.08mJ). These results confirm the importance of high-frequency information in Spiking Transformers: replacing energy-intensive self-attention with lightweight DWC in early stages actually produces better performance.
Table 4: Ablation of Patch-Embedding/ Token-Mixing Strategies on CIFAR100 and CIFAR10‐DVS.
CIFAR100 CIFAR10‐DVS
@@ 4.3 Ablation Study
Overall, Max-Former’s performance gains stem from effective high-frequency information propagation, independent of parameter count, which is evidenced by: (1) Max-pooling patch embedding (Embed-Max/Embed-Max+) consistently outperforms original versions despite similar parameter budgets. (2) Larger DWC kernels (DWC-5/DWC-7) increasing parameters but degrading accuracy (-0.63% on CIFAR100, -2.1% on CIFAR10-DVS vs Max-Former with DWC-3). See Appendix D for visualizations and E for limitations discussion.
4.4 Generality across Convolutional Architectures
We further extend the effectiveness of Max-Former to convolutional architectures by proposing Max-ResNet. The key modification in our Max-ResNet lies in the inclusion of only two additional max-pooling operations compared to MS-ResNet [hu2021advancing]. The detailed implementation of Max-ResNet is provided in Appendix A.3. Training settings are listed in Appendix A.
High-frequency information is essential for SNNs. As shown in Table 5, Max-ResNet achieves a remarkable performance improvement over the baseline MS-ResNet, despite having identical model sizes. Specifically, with the block configuration [2, 2, 2, 2], Max-ResNet improves the CIFAR-10 accuracy by +2.41% (from 94.4% to 96.81%) and the CIFAR-100 accuracy by +6.48%. Similarly, under the [3, 3, 2] configuration, the accuracy increases by +2.25% and +6.65% on CIFAR-10 and CIFAR-100, respectively.
In short, Max‑ResNet‑18 sets state‑of‑the‑art benchmarks across convolutional baselines, with only a moderate model size and an extremely straightforward high-frequency restoration strategy. Therefore, preserving high‑frequency information is fundamental to effective feature representation in SNNs, regardless of architecture.
Table 5: Comparison of different ResNet architectures on CIFAR-10 and CIFAR-100.
Architecture Training Method Block Config. Params (M) Time Step CIFAR-10 Acc. (%) CIFAR-100 Acc. (%)
KDSNN-ResNet-18 [xu2023constructing] Knowledge Distillation [2, 2, 2, 2] 11.22 4 95.72 78.46
MS-ResNet-18 [hu2021advancing] Direct Training [2, 2, 2, 2] 11.22 4 94.40 75.06
[3, 3, 2] 12.50 4 94.92 76.41
MS-ResNet-34 [hu2021advancing] [2, 2, 2, 2] 21.33 4 94.69 75.34
Max-ResNet-18 Direct Training [2, 2, 2, 2] 11.22 4 96.81 (+2.41) 81.54 (+6.48)
[3, 3, 2] 12.50 4 97.17 (+2.25) 83.06 (+6.65)
@@ 5 Conclusion
This work challenges the prevailing assumption that binary activation constraints are the primary cause of SNNs’ performance gap. Through theoretical analysis and empirical validation, we demonstrate for the first time that spiking neurons inherently function as low-pass filters at the network level, resulting in the rapid attenuation of high-frequency components that critically degrade feature representation. We demonstrate that high‑frequency information is crucial for effective spiking computation: Max‑Former (63.99M parameters) achieves 82.39% top‑1 accuracy on ImageNet, outperforming Spikformer (74.81%, 66.34M parameters) by +7.58%, while reducing energy consumption by 30% at a comparable model size; Max‑ResNet‑18 further achieves state‑of‑the‑art performance among convolutional baselines – 97.17% on CIFAR‑10 and 83.06% on CIFAR‑100. Notably, all these improvements are achieved through extremely simple modifications, even slightly reducing the overall model size. We believe this simple yet effective solution will motivate future research to explore the unique properties of SNNs, beyond the established practices in ANN studies.
6 Acknowledgements
This work is supported by the Major Science and Technology Innovation 2030 “Brain Science and Brain-like Research” key project(No.2021ZD0201405), the Guangzhou-HKUST(GZ) Joint Funding Program (Grant No. 2023A03J0682), the National Natural Science Foundation of China (Grant No. 62405255), GuangDong Basic and Applied Basic Research Foundation (No. 2023A1515110679), and partially supported by the collaborative project with Brain Mind Innovation, Inc
Appendix A Analysis of High-Frequency Information in Spiking Transformers
To validate our approach, we conduct experiments across three static and two neuromorphic datasets. In this section, we first provide the detailed experimental setup used to obtain the results presented in our paper. We then perform additional analysis on the importance of high-frequency information in Spiking Transformers. For complete parameter configurations, please refer to our public code repository at https://github.com/bic-L/MaxFormer.
@@ Appendix A Analysis of High-Frequency Information in Spiking Transformers
Resolution $224\times 224$ $\times 32 32×32$$\times 32$ $128\times 128$
@@
A.1 Experimental Details
Datasets: We evaluate Max-Former through comprehensive experiments on static datasets (CIFAR-10 [lecunGradientbasedLearningApplied1998a], CIFAR-100 [krizhevskyLearningMultipleLayers2009a] and ImageNet [dengImagenetLargescaleHierarchical2009]) and neuromorphic datasets (CIFAR10-DVS [liCIFAR10DVSEventStreamDataset2017], DVS128 Gesture [amir2017low]). The training and inference pipeline are implemented in SpikingJelly [SpikingJelly].
Static Datasets: For static image classification, we evaluate on three standard benchmarks. ImageNet-1k [dengImagenetLargescaleHierarchical2009] is one of the most widely used datasets in computer vision. It contains 1.28 million training, 50,000 validation, and 100,000 test images covering the common 1K classes. Both CIFAR-10 [lecunGradientbasedLearningApplied1998a] and CIFAR-100 [krizhevskyLearningMultipleLayers2009a] include 50,000 training images and 10,000 testing images with $\times32$ resolution. The main difference between them is that CIFAR-10 has 10 categories for classification, while CIFAR-100 has 100 categories.
Neuromorphic Datasets: For event-based vision tasks, we evaluate on two standard benchmarks. CIFAR10-DVS [liCIFAR10DVSEventStreamDataset2017] is an event-based version of the CIFAR-10 dataset, created by capturing moving image samples using the Dynamic Vision Sensor (DVS). It includes 10,000 event-based images ($\times128$ pixels) spread across 10 classes, with 9,000 samples for training and 1,000 for testing. The DVS128 Gesture dataset [amir2017low] contains 1,342 event-based recordings of 11 different hand gesture types performed by 29 people under 3 different lighting conditions. Each gesture recording lasts about 6 seconds on average.
Hyper Parameters: Our training scheme mainly follows [zhou2024qkformer] and  [yao2023spike]. Specifically, MixUp [zhang2018mixup], CutMix [yun2019cutmix] and RandAugment [cubuk2020randaugment] are used for data augmentation. The models are trained using AdamW optimizer [loshchilov2017decoupled] with the weight decay of 0.05 for ImageNet-1K classification tasks and the weight decay of 0.06 for all other datasets. Label Smoothing [szegedy2016rethinking] is set as 0.1. Detailed training hyperparameters are shown in Table 6. For our ImageNet experiments, we used 8 NVIDIA A30 GPUs to train most models. However, for the MaxFormer-10-512 (T=4) and MaxFormer-10-768 (T=4) models, we used 8 NVIDIA H20 GPUs instead. For the smaller datasets (CIFAR10, CIFAR100, DVS128 Gesture, and CIFAR10-DVS), we used a single A30 GPU for training.
A.2 Impact of High-Frequency Information on Model Performance
Table 7: Patch embedding and token mixing schemes on CIFAR-100. Acc.:Top-1 Accuracy (%). DWC-NN: spiking depth-wise convolution with kernel size $N\times N$. SSA: Spiking Self-Attention.